OpenStudy (anonymous):

how do you solve x(x^2-x-2), [-1,1] using the mean value theorem

OpenStudy (anonymous):

this question is ill posed. do you mean "what does the mean value theorem say about the function \[f(x)=x(x^2-x-2)\] over the interval [-1,1]?

OpenStudy (amistre64):

\[\frac{f(1)-f(-1)}{1--1}=f'(c)\]

OpenStudy (anonymous):

or does it mean "find the value of c so that \[f'(c)=\frac{f(1)-f(-1)}{2}\]

OpenStudy (amistre64):

i assume of course the second :) but yeah, it could mean the slope obtained ....